# Learning Spatial-Aware Manipulation Ordering

### v1

@@ 1 Introduction
Object manipulation [1, 2, 3] in real-world environments is a fundamental capability for robots to interact with diverse objects to perform tasks [4]. In cluttered environments, objects are often closely packed, partially occluded, or physically constrained by one another [5]. The order in which objects are manipulated can significantly impact both task efficiency and scene stability. Improper ordering may lead to collisions or structural collapse [6]. Effective manipulation ordering requires understanding how the spatial layout of objects constrains their interactions. Capturing spatial conditions and inferring manipulation order is essential for enabling safe robotic manipulation.
Most existing approaches for manipulation in cluttered environments do not explicitly model manipulation ordering. As shown in Fig. 1(a), some systems [7, 8, 9] systems apply heuristic algorithm after object recognition. These pipelines are generally limited in their ability to generalize and tend to break down when spatial relations vary across different scenes. More recent methods, shown in Fig. 1(b), leverage large vision-language models to guide the manipulation sequence through explicit prompts. While these approaches are flexible, they suffer from long inference times [10], often spanning several seconds, making them unsuitable for real-time deployment. These limitations motivate the need for a unified framework that can directly learn spatial-aware manipulation ordering with low latency and high adaptability.
@@
As illustrated in Fig. 1(c), we propose OrderMind, a unified framework that learns spatial-aware manipulation priorities directly from cluttered environments. It combines a spatial context understanding module and a temporal priority structuring module to infer manipulation ordering based on local geometry and interactions. A k-Nearest Neighbors algorithm-based spatial graph is proposed to capture object-object and object-manipulator relations, supporting efficient spatial layout understanding. We introduce a spatial prior labeling method that guides a vision-language model to generate plausible orderings. This enables training without manual labels. We modify a vision-centric backbone [11] for spatial context understanding. By integrating object understanding and order prediction into a single inference pass, the model achieves accurate and efficient manipulation ordering, maintaining spatial consistency across diverse cluttered scenes while supporting real-time performance.
@@ 2.1 Models for Robotic Manipulation
With diverse modeling paradigms ranging from vision-language reasoning to physics-based dynamics, recent models for robotic manipulation aim to bridge perception and control through both learned and analytical representations. Some vision-language models, such as CLIPort [12] and PaLM-E [13], extend manipulation capabilities by conditioning on language and integrating multimodal information from vision and proprioception. SayCan [14] introduced affordance-aware language grounding. RT-1 [15] and its successor RT-2 [16] re-frame actions as text tokens and co-finetune on internet-scale vision-language corpora, unlocking chain-of-thought reasoning for semantic and arithmetic object commands. VIMA [17] generalises prompt engineering to robotics for strong zero-shot performance.
ReKep [18] leveraged the power of VLM, specifically GPT-4o [19], to effectively reason about 3D rotations. It achieved this by expressing spatial relationships as arithmetic operations between keypoints. Moreover, advances in perception method provide scene reconstruction [20] and object understanding [21, 22] that can be used to preprocess visual inputs before manipulation. SeeDo [23] utilizes VLM with chain-of-thought prompting and visual cues from video tracking to enable effective reasoning for robotic task planning. SpatialVLM [24] enhanced VLM spatial reasoning capabilities by training on a large-scale generated spatial Visual Question Answering dataset. Such perception-driven priors enhance spatial awareness [25, 26] and improve the reliability of downstream control. RoboMamba [27] integrated a vision encoder with Mamba [28] Large Language Model to obtain both visual common sense and robotic-related reasoning abilities. Although VLM-based models possess general reasoning capability, they are relatively slow and cannot meet the real-time requirements of robotic manipulation tasks.
@@ 2.2 Manipulation Relationship Network
Early methods in visual relationship detection focused on building subject-predicate-object triplets and combining linguistic priors or global context for joint learning. Framework [29] learned the appearance of objects and predicates individually and used the relationship embedding space learned from language to detect visual relationships. SCG [30] applied interactive message passing for deep object-relation interactions. VRL [31] used a reinforcement learning-based method to explore multiple relationships and attributes sequentially. CDDN Network [32] utilized word semantics and visual scene graphs to capture the interactions between different object instances for more accurate predictions. Graph R-CNN [33] leverages object-relationship regularities to generate scene graphs from detected objects in an end-to-end manner. Motifnet [34] staged detection results and relationships to establish a rich context for subsequent prediction.
To further integrate the physical order of actions into robotic manipulation. VMRN [35] proposed a novel pooling layer to predict manipulation relations in parallel. GVMRN [36] adopted GCNs to enhance contextual reasoning and filter irrelevant pairs, boosting inference robustness. MRG [37] used affordance representation to get the potential manipulation relationships of an arbitrary scene. VeriGraph[38] generated high-level task plans for robot object manipulation based on the scene graph. D3G[39] detected objects and reasoned relationships for grasp planning simultaneously by leveraging a transformer-based detector and graph layers. Although these models are capable of generating object relation graphs that can be used for manipulation tasks, they still focus on perceiving relationships between objects, ignoring the relationships between objects and the robot, and still require further post-processing to obtain the manipulation order.
@@ 3.2 Spatial-Aware Manipulation Ordering
Spatial Context Understanding. Occlusions and inter-object support in cluttered scenes make it necessary to reason about spatial and physical relationships rather than visual appearance alone. To enable accurate manipulation ordering in cluttered environments, we focus on extracting spatial-aware representations that capture both local and global scene structure. We represent each object using its 3D bounding box center and associated intrinsic and extrinsic attributes. The intrinsic attributes describe inherent properties such as semantic class and physical extent, while the extrinsic ones define the object’s pose in the world coordinate system. The object centers collectively form a sparse point cloud representing the 3D scene layout. Based on this point cloud, we construct a spatial graph where nodes correspond to objects and edges reflect geometric proximity.
To model inter-object spatial dependencies, we apply a k-Nearest Neighbors strategy to build a localized spatial graph around each object. As shown in Fig. 3, we treat the center point of each object as a vertex in the graph and connect each object to its neighbors with an edge. Each vertex in this graph incorporates both intrinsic and extrinsic attributes. For each center $p_{j}$, messages from its spatial neighbors $\mathcal{N}_{k}({i})$ are aggregated to form spatial-aware features that encode local geometric structure, as described in Equation 1:
$\operatorname{Fusion}\left({{f_{i},f_{j}}}\right)=\mathcal{M}\!\left(\operatorname{Linear}\!\left(\operatorname{Concate}\!\left(f_{i},f_{j}-f_{i}\right)\right)\right),~~~~\forall{p_{j}}\in\mathcal{N}_{k}({i}),$ (1)
This design captures how physical proximity and spatial alignment influence the feasibility of manipulation actions. In line with PointNet-style architectures [40, 41], feature aggregation is performed using max pooling $\mathcal{M}$ across neighborhoods to produce compact object-level embeddings.
@@
Firstly, a set of object ordering tokens extracted from the image encoder is aggregated via global max pooling to form a high-level scene representation, denoted as $G$. This global context encodes the layout of the scene, capturing occlusion patterns, object stacking, and spatial symmetry. Each object token $Q$ is then refined through self-attention, which incorporates inter-object interactions and aligns the object tokens with one another. Formally, the attention flow is defined as:
@@
Preferential Order Alignment. To model spatial-aware manipulation ordering from supervision in the form of discrete annotated orders, we first align predicted and ground-truth objects via bipartite matching. Let $o$ and $\hat{o}$ denote ground-truth and predicted objects, respectively. We formulate the assignment using the Hungarian algorithm, where object similarity serves as the matching cost $\hat{\sigma}=\mathop{\arg\min}\limits_{\sigma}\sum\limits_{i=1}^{N}\mathcal{L}_{\text{hungarian}}\left(o_{i},\hat{o}_{\sigma(i)}\right)$. When the object correspondence $\hat{\sigma}$ established, we align the prediction of ordering preference with ground truth order by comparing pairwise relations. Instead of directly predicting discrete ranks, the model learns to produce continuous scores $\hat{s}$ to implicitly align sorted order with the annotated order. The objective is expressed as follows:
@@ 3.3 Spatial Priors Order Labeling
Learning spatial-aware manipulation ordering relies on well-structured annotations that reflect object interaction priorities in cluttered environments. In our framework, these manipulation orders are generated by a VLM [42]. To enhance the quality and consistency of these labels, we introduce two spatial priors to serve as auxiliary signals during the generation process. These priors influence how the large model interprets spatial relationships, ensuring that the predicted orderings align with common physical and operational patterns observed in cluttered scenes.
@@
Local Optimality Prior. This prior captures vertical accessibility by identifying objects that are not occluded from above. In cluttered environments, vertical stacking is frequent, and selecting objects from the top layers helps avoid disturbing other items and maintains the scene’s stability. For object $\mathbf{o}_{i}\in\mathbf{O}$ and its 3D bounding box $B_{i}$, we define $z_{\max}(\mathbf{o}_{i})=\sup\left\{z\mid(x,y,z)\in B_{i}\right\}$ as the height of its topmost point. The volume above $\mathbf{o}_{i}$ is defined as the upward extension of its projected area:
@@ 4.1.1 Environments and Task
We constructed datasets involving complex robotic manipulation scenarios in both simulated and real-world environments. For simulation, we utilized Pybullet [43] as the engine and designed cluttered scenes using the YCB object set [44]. Three levels of difficulty were established: Easy with 24 objects, Medium with 36 objects, and Hard with 60 objects. The simulation dataset consists of 161,722 RGB-D images for training, 1,500 images for validation. Closed-loop testings are performed within the PyBullet simulator. For real-world experiments, we collected 26,324 images for training and 6,581 for validation. Closed-loop testings are directly conducted in real-world cluttered environments.
@@ 4.1.2 Training and Implementation Details
We trained our model on a single RTX 4090 GPU using a batch size of 24. The optimizer is AdamW [45] with an initial learning rate of $2\times 10^{-4}$. The learning rate follows a cosine annealing schedule [46] with a final minimum value of 0.01. Training proceeds in two stages. The first stage focuses on pre-training the image backbone [47, 48] for 50 epochs. In the second stage, the full model is trained end-to-end for 10 epochs, enabling the system to learn order-aware manipulation directly from sufficient spatial representation.
@@ 4.2.1 Simulation Experiments
Tab. 1 presents a comprehensive evaluation of manipulation performance in cluttered environments across four experimental settings: heuristic pipeline, two-stage framework, privilege information-based two-stage framework, and our proposed unified spatial-aware framework. Privilege information refers to models having direct access to ground truth object pose information in the simulation environment. We derive our Spatial Priors (Sec. 3.3) as the Spatial Priors Heuristic (SPH) for complete comparison. The speed of VLMs is calculated by using OpenRouter [55].
We first evaluate the upper bound of vision-language reasoning using VLMs with privileged information. Even with privileged information, the best model achieves only 92.5% success rate in easy setting and 78.5% in hard setting. Next, we test heuristic pipelines using YOLOv11 [53] and UniDet3D [54], selecting objects by Spatial Priors Heuristic, confidence, and distance to the end effector. Although UniDet3D and YOLO has fewer parameter counts, it is limited to 0.8 FPS due to its point cloud-based architecture. We also evaluate two-stage frameworks where UniDet3D results are passed to VLMs for ordering, which achieves only 0.1 FPS and cannot meet the real-time demand.
@@ Appendix B Data Annotation Pipeline
This section details our methodology for generating spatial layout perception labels within real-world environments. To elaborate, for a given RGB image, we initially employ Segment Anything [56] to produce instance-level segmentation masks. These masks are subsequently integrated with corresponding RGB-D data to create instance-level point clouds. We employ the depth inpainting method [57] to reduce noise during the data annotation stage. Following this, the RANSAC algorithm [58] is utilized to extract the dominant planar surface from these point clouds.
@@ C.1 Comparison with Baselines under Multiple Difficulty Mode
These tables demonstrate that our OrderMind method achieves the best performance across easy, moderate, and hard difficulty levels. It is also superior in performance and faster than the VLM method with privilege information. Notably, OrderMind’s performance advantage becomes more pronounced as the difficulty increases. Importantly, ’fail num’ indicates the number of failed attempts. Thus, inaccurate detection results can lead the robotic end-effector to repeatedly try to pick false-positive objects, resulting in a higher num of failure attempts. So that experiments relying on UniDet3D [54] usually have a higher ’fail num’ metric. A lower residual count, lower object disturbance, and higher success rate demonstrate OrderMind’s strong understanding of scene stability and its ability to perform object manipulation quickly and accurately.
@@
GPT-4o [19] 19.2 2.1 4.8 2.3 90.3
Claude-3.7-Sonnet [49] 23.3 2.1 0.7 2.0 92.1
Gemini-2.5 [50] 23.3 2.1 0.7 1.6 92.4
Gemma-3 [51] 22.8 3.4 1.2 2.2 88.3
InternVL3 [52] 22.6 3.3 1.4 2.3 88.2
Qwen2.5-VL [42] 23.1 2.1 0.9 1.8 92.5
UniDet3D [54]+Confidence 20.9 26.4 3.1 2.5 49.4
@@
UniDet3D+InternVL3 21.2 36.0 2.8 2.9 39.3
UniDet3D+Qwen2.5-VL 20.5 39.0 3.5 3.4 39.1

### v2

@@ 1 Introduction
Object manipulation [ehsani2021manipulathor, yin2021modeling, shridhar2023perceiver] in real-world environments is a fundamental capability for robots to interact with diverse objects to perform tasks [eom2024mogrip]. In cluttered environments, objects are often closely packed, partially occluded, or physically constrained by one another [danielczuk2021object]. The order in which objects are manipulated can significantly impact both task efficiency and scene stability. Improper ordering may lead to collisions or structural collapse [xu2023optimal]. Effective manipulation ordering requires understanding how the spatial layout of objects constrains their interactions. Capturing spatial conditions and inferring manipulation order is essential for enabling safe robotic manipulation.
Most existing approaches for manipulation in cluttered environments do not explicitly model manipulation ordering. As shown in Fig. 1(a), some systems [saxena2023planning, haustein2019object, KARAMI2025104943] systems apply heuristic algorithm after object recognition. These pipelines are generally limited in their ability to generalize and tend to break down when spatial relations vary across different scenes. More recent methods, shown in Fig. 1(b), leverage large vision-language models to guide the manipulation sequence through explicit prompts. While these approaches are flexible, they suffer from long inference times [kapelyukh2024dream2real], often spanning several seconds, making them unsuitable for real-time deployment. These limitations motivate the need for a unified framework that can directly learn spatial-aware manipulation ordering with low latency and high adaptability.
@@
As illustrated in Fig. 1(c), we propose OrderMind, a unified framework that learns spatial-aware manipulation priorities directly from cluttered environments. It combines a spatial context understanding module and a temporal priority structuring module to infer manipulation ordering based on local geometry and interactions. A k-Nearest Neighbors algorithm-based spatial graph is proposed to capture object-object and object-manipulator relations, supporting efficient spatial layout understanding. We introduce a spatial prior labeling method that guides a vision-language model to generate plausible orderings. This enables training without manual labels. We modify a vision-centric backbone [zhou2019objects] for spatial context understanding. By integrating object understanding and order prediction into a single inference pass, the model achieves accurate and efficient manipulation ordering, maintaining spatial consistency across diverse cluttered scenes while supporting real-time performance.
@@ 2.1 Models for Robotic Manipulation
With diverse modeling paradigms ranging from vision-language reasoning to physics-based dynamics, recent models for robotic manipulation aim to bridge perception and control through both learned and analytical representations. Some vision-language models, such as CLIPort [shridhar2022cliport] and PaLM-E [driess2023palm], extend manipulation capabilities by conditioning on language and integrating multimodal information from vision and proprioception. SayCan [ahn2022can] introduced affordance-aware language grounding. RT-1 [brohan2022rt] and its successor RT-2 [brohan2023rt] re-frame actions as text tokens and co-finetune on internet-scale vision-language corpora, unlocking chain-of-thought reasoning for semantic and arithmetic object commands. VIMA [jiang2022vima] generalises prompt engineering to robotics for strong zero-shot performance.
ReKep [huang2024rekep] leveraged the power of VLM, specifically GPT-4o [achiam2023gpt], to effectively reason about 3D rotations. It achieved this by expressing spatial relationships as arithmetic operations between keypoints. Moreover, advances in perception method provide scene reconstruction [yan2024pointssc] and object understanding [guo2025dark, gaodeep] that can be used to preprocess visual inputs before manipulation. SeeDo [wang2024vlm] utilizes VLM with chain-of-thought prompting and visual cues from video tracking to enable effective reasoning for robotic task planning. SpatialVLM [chen2024spatialvlm] enhanced VLM spatial reasoning capabilities by training on a large-scale generated spatial Visual Question Answering dataset. Such perception-driven priors enhance spatial awareness [li2025constrained, li2025papl] and improve the reliability of downstream control. RoboMamba [liu2024robomamba] integrated a vision encoder with Mamba [gu2023mamba] and Large Language Model to obtain both visual common sense and robotic-related reasoning abilities. Although VLM-based models possess general reasoning capability, they are relatively slow and cannot meet the real-time requirements of robotic manipulation tasks.
@@ 2.2 Manipulation Relationship Network
Early methods in visual relationship detection focused on building subject-predicate-object triplets and combining linguistic priors or global context for joint learning. Visual relationship prediction model [lu2016visual] learned the appearance of objects and predicates individually and used the relationship embedding space learned from language to detect visual relationships. SCG [xu2017scene] applied interactive message passing for deep object-relation interactions. VRL [liang2017deep] used a reinforcement learning-based method to explore multiple relationships and attributes sequentially. CDDN Network [cui2018context] utilized word semantics and visual scene graphs to capture the interactions between different object instances for more accurate predictions. Graph R-CNN [yang2018graph] leverages object-relationship regularities to generate scene graphs from detected objects in an end-to-end manner. Motifnet [zellers2018neural] staged detection results and relationships to establish a rich context for subsequent prediction.
To further integrate the physical order of actions into robotic manipulation. VMRN [zhang2018visual] proposed a novel pooling layer to predict manipulation relations in parallel. GVMRN [zuo2021graph] adopted GCNs to enhance contextual reasoning and filter irrelevant pairs, boosting inference robustness. MRG [tang2021relationship] used affordance representation to get the potential manipulation relationships of an arbitrary scene. VeriGraph[ekpo2024verigraph] generated high-level task plans for robot object manipulation based on the scene graph. D3G[rabino2025modern] detected objects and reasoned relationships for grasp planning simultaneously by leveraging a transformer-based detector and graph layers. Although these models are capable of generating object relation graphs that can be used for manipulation tasks, they still focus on perceiving relationships between objects, ignoring the relationships between objects and the robot, and still require further post-processing to obtain the manipulation order.
@@ 3.2 Spatial-Aware Manipulation Ordering
Spatial Context Understanding. Occlusions and inter-object support in cluttered scenes make it necessary to reason about spatial and physical relationships rather than visual appearance alone. To enable accurate manipulation ordering in cluttered environments, we focus on extracting spatial-aware representations that capture both local and global scene structure. We represent each object using its 3D bounding box center and associated intrinsic and extrinsic attributes. The intrinsic attributes describe inherent properties such as semantic class and physical extent, while the extrinsic ones define the object’s pose in the world coordinate system. The object centers collectively form a sparse point cloud representing the 3D scene layout. Based on this point cloud, we construct a spatial graph where nodes correspond to objects and edges encode geometric proximity.
To model inter-object spatial dependencies, we apply a k-Nearest Neighbors strategy to build a localized spatial graph around each object. As shown in Fig. 3, we treat the center point of each object as a node in the graph and connect each object to its neighbors with edges. Each node in this graph incorporates both intrinsic and extrinsic attributes. For each center $p_{i}$, messages from its spatial neighbors $\mathcal{N}_{k}({p_{i}})$ are aggregated to form spatial-aware features that encode local geometric structure, as described in Equation 1:
$\operatorname{Fusion}\left({{f_{i},f_{j}}}\right)=\mathcal{M}\!\left(\operatorname{Linear}\!\left(\operatorname{Concate}\!\left(f_{i},f_{j}-f_{i}\right)\right)\right),~~~~\forall{p_{j}}\in\mathcal{N}_{k}({p_{i}}),$ (1)
This design captures how physical proximity and spatial alignment influence the feasibility of manipulation actions. In line with PointNet-style architectures [qi2017pointnet, yu2023adapointr], feature aggregation is performed using max pooling $\mathcal{M}$ across neighborhoods to produce compact object-level embeddings.
@@
Firstly, a set of object ordering tokens extracted from the image encoder is aggregated via a global max pooling to form a high-level scene representation, denoted as $G$. This global context encodes the layout of the scene, capturing occlusion patterns, object stacking, and spatial symmetry. Each object token $Q$ is then refined through self-attention, which incorporates inter-object interactions and aligns the object tokens with one another. Formally, the attention flow is defined as:
@@
Preferential Order Alignment. To learn a spatially-aware manipulation ordering from discrete order annotations, we first align predicted and ground-truth objects via bipartite matching. Let $o$ and $\hat{o}$ denote ground-truth and predicted objects, respectively. We formulate the assignment using the Hungarian algorithm, where object similarity serves as the matching cost $\hat{\sigma}=\mathop{\arg\min}\limits_{\sigma}\sum\limits_{i=1}^{N}\mathcal{L}_{\text{hungarian}}\left(o_{i},\hat{o}_{\sigma(i)}\right)$. When the object correspondence $\hat{\sigma}$ established, we align the prediction of ordering preference with ground truth order by comparing pairwise relations. Instead of directly predicting discrete ranks, the model learns to produce continuous scores $\hat{s}$ to implicitly align sorted order with the annotated order. The objective is expressed as follows:
@@ 3.3 Spatial Priors Order Labeling
Learning spatial-aware manipulation ordering relies on well-structured annotations that reflect object interaction priorities in cluttered environments. In our framework, these manipulation orders are generated by a VLM [Qwen2.5-VL]. To enhance the quality and consistency of these labels, we introduce two spatial priors to serve as auxiliary signals during the generation process. These priors influence how the large model interprets spatial relationships, ensuring that the predicted orderings align with common physical and operational patterns observed in cluttered scenes.
@@
Local Optimality Prior. This prior captures vertical accessibility by identifying objects that are not occluded from above. In cluttered environments, vertical stacking is frequent, and selecting objects from the top layers helps avoid disturbing other items and maintains the scene’s stability. For object $\mathbf{o}_{i}\in\mathbf{O}$ and its 3D bounding box $B_{i}$, we define $z_{\max}(\mathbf{o}_{i})=\sup\left\{z\mid(x,y,z)\in B_{i}\right\}$ as the height of its topmost point. The volume above $\mathbf{o}_{i}$ is defined as the upward extension of its projected area $A_{i}$:
@@ 4.1.1 Environments and Task
We constructed datasets involving complex robotic manipulation scenarios in both simulated and real-world environments. For simulation, we utilized Pybullet [coumans2021] as the engine and designed cluttered scenes using the YCB object set [7251504]. Three levels of difficulty were established: Easy with 24 objects, Medium with 36 objects, and Hard with 60 objects. The simulation dataset consists of 161,722 RGB-D images for training, 1,500 images for validation. Closed-loop testings are performed within the PyBullet simulator. For real-world experiments, we collected 26,324 images for training and 6,581 for validation. Closed-loop testings are directly conducted in real-world cluttered environments.
@@ 4.1.2 Training and Implementation Details
We trained our model on a single RTX 4090 GPU using a batch size of 24. The optimizer is AdamW [loshchilovdecoupled] with an initial learning rate of $2\times 10^{-4}$ and a weight decay of 0.01. The learning rate follows a cosine annealing schedule [loshchilov2022sgdr]. Training proceeds in two stages. The first stage focuses on pre-training the image backbone [yu2018deep, zhou2020tracking] for 50 epochs. In the second stage, the full model is trained end-to-end for 10 epochs, enabling the system to learn order-aware manipulation directly from sufficient spatial representation.
@@ 4.2.1 Simulation Experiments
Tab. 1 presents a comprehensive evaluation of manipulation performance in cluttered environments across four experimental settings: heuristic pipeline, two-stage framework, privilege information-based two-stage framework, and our proposed unified spatial-aware framework. Privilege information refers to models having direct access to ground truth object pose information in the simulation environment. We derive our Spatial Priors (Sec. 3.3) as the Spatial Priors Heuristic (SPH) for complete comparison. The speed of VLMs is calculated by using OpenRouter [openrouter].
We first evaluate the upper bound of vision-language reasoning using VLMs with privileged information. Even with privileged information, the best model achieves only 92.5% success rate in easy setting and 78.5% in hard setting. Next, we test heuristic pipelines using YOLOv11 [ultralytics2025] and UniDet3D [kolodiazhnyi2025unidet3d], selecting objects by Spatial Priors Heuristic, confidence, and distance to the end effector. Although UniDet3D and YOLO has fewer parameter counts, it is limited to 0.8 FPS due to its point cloud-based architecture. We also evaluate two-stage frameworks where UniDet3D results are passed to VLMs for ordering, which achieves only 0.1 FPS and cannot meet the real-time demand.
@@ Appendix B Data Annotation Pipeline
This section details our methodology for generating spatial layout perception labels within real-world environments. To elaborate, for a given RGB image, we initially employ Segment Anything [kirillov2023segment] to produce instance-level segmentation masks. These masks are subsequently integrated with corresponding RGB-D data to create instance-level point clouds. We employ the depth inpainting method [Silberman:ECCV12] to reduce noise during the data annotation stage. Following this, the RANSAC algorithm [fischler1981random] is utilized to extract the dominant planar surface from these point clouds.
@@ C.1 Comparison with Baselines under Multiple Difficulty Mode
These tables demonstrate that our OrderMind method achieves the best performance across easy, moderate, and hard difficulty levels. It is also superior in performance and faster than the VLM method with privilege information. Notably, OrderMind’s performance advantage becomes more pronounced as the difficulty increases. Importantly, ’fail num’ indicates the number of failed attempts. Thus, inaccurate detection results can lead the robotic end-effector to repeatedly try to pick false-positive objects, resulting in a higher num of failure attempts. So that experiments relying on UniDet3D [kolodiazhnyi2025unidet3d] usually have a higher ’fail num’ metric. A lower residual count, lower object disturbance, and higher success rate demonstrate OrderMind’s strong understanding of scene stability and its ability to perform object manipulation quickly and accurately.
@@
GPT-4o [achiam2023gpt] 19.2 2.1 4.8 2.3 90.3
Claude-3.7-Sonnet [claude3.7] 23.3 2.1 0.7 2.0 92.1
Gemini-2.5 [google2025gemini2.5] 23.3 2.1 0.7 1.6 92.4
Gemma-3 [team2025gemma] 22.8 3.4 1.2 2.2 88.3
InternVL3 [zhu2025internvl3] 22.6 3.3 1.4 2.3 88.2
Qwen2.5-VL [Qwen2.5-VL] 23.1 2.1 0.9 1.8 92.5
UniDet3D [kolodiazhnyi2025unidet3d]+Confidence 20.9 26.4 3.1 2.5 49.4
@@
UniDet3D+InternVL3 21.2 36.0 2.8 2.9 38.3
UniDet3D+Qwen2.5-VL 20.5 39.0 3.5 3.4 38.8